OpenStudy (anonymous):

Use the distributive property to simplify the expression. –7(8x – 3) + x A. x + 10 B. –15x + 4 C. –9x – 4 D. –55x + 21

OpenStudy (anonymous):

-7 * 8 = -56, so we get -56x. -7 * -3 = +21

OpenStudy (anonymous):

-56x + 21 + x = -55x + 21

OpenStudy (mathstudent55):

Here is a similar example. Follow along and then try your problem. -2(4x + 3) + x = (-2)(4x) + (-2)(3) + x = -8x - 6 + x = -7x + 6

OpenStudy (mrnood):

@AleshGames it is prohibited in the code of conduct to give complete direct answers. You will help the OP MORE by guiding him to make his OWN calculation